Find the APR and ieff
Payday Loan
Interest is $20 on a $300 loan for a 2 week period.

My teacher put 20/300=6.7% for the 2 weeks
And then 6.7x26 = 173 APR
173 is the correct anwser,
Where does the 26 come from?
Answer #1

Hello,

Please find the answer below:

APR stands for Annualized Percentage Rate.

Here the interest mentioned is for a 2 week period.

So for a 2 week period , Interest = ?

=Interest amount/loan amount = 20/300 = 6.7%

Now we need to convert this 2 week period interest rate in an annual rate

1 year = 52 weeks, so time period as per 2 week period = 52/2 = 26 weeks.

Interest rate for 1 year = 26 * 6.7 = 173 (its actually 174) Annual Percentage Rate

I hope this helps you understand the calculation of 26 in the solution.

Thanks